### v5

# Discriminative and Consistent Representation Distillation

Abstract
Knowledge Distillation (KD) aims to transfer knowledge from a large teacher model to a smaller student model. While contrastive learning has shown promise in self-supervised learning by creating discriminative representations, its application in knowledge distillation remains limited and focuses primarily on discrimination, neglecting the structural relationships captured by the teacher model. To address this limitation, we propose Discriminative and Consistent Distillation (DCD), which employs a contrastive loss along with a consistency regularization to minimize the discrepancy between the distributions of teacher and student representations. Our method introduces learnable temperature and bias parameters that adapt during training to balance these complementary objectives, replacing the fixed hyperparameters commonly used in contrastive learning approaches. Through extensive experiments on CIFAR-100 and ImageNet ILSVRC-2012, we demonstrate that DCD achieves state-of-the-art performance, with the student model sometimes surpassing the teacher’s accuracy. Furthermore, we show that DCD’s learned representations exhibit superior cross-dataset generalization when transferred to Tiny ImageNet and STL-10111Code is available at: https://github.com/giakoumoglou/distillers..
Figure 1: Overview of DCD. (a) Discriminative learning through contrastive distillation encourages student features (solid blue) to differentiate between instances by pulling them closer to their corresponding teacher features (transparent blue) while pushing away from other instances as negative samples (black dots). (b) Structural consistency through consistency regularization preserves the distributional relationship patterns captured by the teacher model by aligning the student and teacher feature similarities (represented by dotted lines) through KL divergence minimization.
1 Introduction
Knowledge Distillation (KD) has emerged as a prominent technique for model compression, enabling the transfer of knowledge from large, high-capacity teacher models to more compact student models [33]. This approach is particularly relevant today, as state-of-the-art vision models in tasks such as image classification [48, 21], object detection [60, 44], and semantic segmentation [10, 11] continue to grow in size and complexity. While these large models achieve impressive performance, their computational demands make them impractical for real-world applications [23, 39], leading practitioners to seek more efficient alternatives through model compression techniques [5, 59].
The representation learning capabilities of neural networks play a crucial role in their performance [41, 4]. In the context of KD, while the original approach [33] focused on transferring knowledge through logit outputs, subsequent work has emphasized the importance of intermediate feature representations [61, 74, 67, 56]. These intermediate representations capture rich structural information and hierarchical features that are essential for robust model performance [38, 77]. Recent advances in KD have explored various ways to transfer this representational knowledge, including attention transfer [74], correlation congruence [58], and relational knowledge distillation [56].
Contrastive learning has recently revolutionized self-supervised representation learning [13, 30], demonstrating its effectiveness in learning discriminative features without labels. This success has inspired its adoption in KD frameworks [66, 22]. However, existing contrastive distillation approaches face several limitations: they often require large memory banks to store negative samples [66], rely on fixed hyperparameters that limit their adaptability [13], and may not fully preserve the structural relationships captured by the teacher model [67]. Furthermore, the focus on discrimination alone can lead to suboptimal knowledge transfer, as it neglects the importance of maintaining consistent representational patterns between teacher and student models [66, 58].
To address these limitations, we propose Discriminative and Consistent Distillation (DCD), an approach that combines contrastive learning with consistency regularization to ensure both discriminative power and structural consistency in the student’s representations (see Figure 1). Our method eliminates the need for memory banks by leveraging in-batch negative samples and introduces learnable temperature and bias parameters that dynamically adjust during training, enabling more flexible and efficient knowledge transfer. This adaptive approach allows the student to better capture both instance-level discriminative features and global structural patterns from the teacher, leading to more robust and generalizable representations [38].
Our contributions are twofold:
  1. 1.
    We propose an approach that combines contrastive learning with consistency regularization to ensure both discriminative and structurally consistent representations. Our method eliminates the need for memory banks and introduces learnable parameters that dynamically adjust during training, enabling more efficient knowledge transfer than existing approaches.
  2. 2.
    We demonstrate the effectiveness of our approach through extensive experiments on standard benchmarks, showing significant improvements in both accuracy and robustness. DCD outperforms other methods, achieving a 20.31% relative improvement [^2] Average relative improvement is calculated as: $\frac{1}{N}\sum_{i=1}^{N}\frac{\text{Acc}^{i}_{\text{DCD}}-\text{Acc}^{i}_{\text{KD}}}{\text{Acc}^{i}_{\text{KD}}-\text{Acc}^{i}_{\text{van}}}$, where $\text{Acc}^{i}_{\text{DCD}}$, $\text{Acc}^{i}_{\text{KD}}$, and $\text{Acc}^{i}_{\text{van}}$ represent the accuracies of DCD, KD, and vanilla training of the $i$-th student model, respectively [66]. over the original KD. When combined with KD, it shows a 73.87% relative improvement over the original KD.
The rest of this paper is organized as follows. Section 2 reviews related work in knowledge distillation and contrastive learning. Section 3 details our proposed methodology. Section 4 presents our experimental setup and results, and Section 5 concludes the paper.
2 Related Work
Our method combines contrastive learning with knowledge distillation by adding consistency regularization.
2.1 Knowledge Distillation
The original knowledge distillation work by [33] introduced transferring knowledge through softened logit outputs using temperature scaling in the softmax. Similar to [33], our method uses temperature scaling but learns the optimal temperature during training for better knowledge transfer.
Logit-based methods.
Several methods have improved logit-based distillation through techniques like label decoupling [81], instance-specific label smoothing [73], probability reweighting [55], and normalizing logits before softmax and KL divergence [65]. Some works focus on dynamic temperature adjustment [42] and separate learning of response-based and feature-based distillation [79]. Additional improvements include transformations for better teacher-student alignment [80] and methods for knowledge transfer from stronger teachers [35]. Unlike existing methods such as [42] that use fixed temperature schedules and MLPs, our approach introduces truly learnable temperature parameters that adapt during training, making it more flexible and efficient.
Feature-based methods.
The work by [61] used intermediate feature hints to guide student learning, while [74] aligned attention maps between teacher and student. Several methods have explored structural relationships: [58] preserved feature space relationships, [45] ensured functional consistency, and [25] transferred class-level attention information. The method by [56] transferred mutual relations between data examples. Recent works have introduced cross-stage connection paths [12], direct reuse of teacher’s classifier [8], and many-to-one representation matching [47]. Contrastive Representation Distillation (CRD) [66] used contrastive learning to maximize mutual information between representations but needed large memory buffers. Compared to [66], our approach uses efficient in-batch sampling and adds structural consistency, making it more practical. Our method also differs from previous approaches like [35, 58, 57, 67] in how we define and preserve structural relationships.
Architecture-aware methods.
Recent studies have examined how network architecture affects distillation success. Some works developed meta-learning for architecture search [19], training-free student architecture selection [20], and graph-based architecture adaptation [50]. Methods by [28] and [52] addressed distillation between different architectures through unified feature spaces and intermediate networks respectively. These approaches aim to optimize not just the distillation process but also the underlying network structures to achieve better knowledge transfer.
2.2 Contrastive Learning
Contrastive methods in self-supervised learning have proven effective for learning robust representations by maximizing mutual information [34, 68]. These methods transform unsupervised learning into a classification problem, building on foundational work in metric learning [16, 27] to distinguish between positive and negative samples. The theoretical foundations [2, 26] show that such objectives maximize a lower bound on mutual information, crucial for meaningful representations. Recent advances using momentum encoders [30], stronger augmentations [13], and methods that eliminate negative samples [24] have further improved self-supervised learning. Additionally, recent strategies explore invariance regularizers [54], while others prevent model collapse through redundancy reduction [76] or regularization [3]. Some approaches achieve this by eliminating negative samples through asymmetric Siamese structures or normalization [24, 14, 6]. Our method combines instance-level discrimination [70, 71] with an consistency constraint, ensuring the student learns both discriminative features and preserves the teacher’s structural knowledge. Furthermore, our approach does not rely on fixed negative samples or momentum encoders [51]; instead, it employs a dynamic method that adapts to the model’s current state during training. Our method shares the theoretical underpinnings of mutual information maximization but extends this framework to include explicit structural preservation, providing a more comprehensive approach to knowledge transfer.
3 Methodology
This section presents our methodology to improve the efficiency and accuracy of KD. Our method, Discriminative and Consistent Distillation (DCD), focuses on learning representations that are both discriminative through contrastive learning and structurally consistent with the teacher model through a consistency regularization. Our method ensures that the student model learns to differentiate between different instances while preserving the distributional relationships captured by the teacher model. Figure 1 shows an overview of the proposed method in the latent space.
3.1 Preliminaries
KD involves transferring knowledge from a high-capacity teacher neural network, denoted a $f^{T}$, to a more compact student neural network $f^{S}$. Consider $x_{i}$ as the input to these networks, typically an image. We represent the outputs at the penultimate layer (just before the final classification layer, or logits) as $\mathbf{z}_{i}^{T}=f^{T}(x_{i})$ and $\mathbf{z}_{i}^{S}=f^{S}(x_{i})$ for the teacher and student models, respectively. The primary objective of KD is to enable the student model to approximate the performance of the teacher model. The overall distillation process can be mathematically expressed as:
$\mathcal{L}=\mathcal{L}_{\text{sup}}(y_{i},\mathbf{z}_{i}^{S})+\lambda\cdot\mathcal{L}_{\text{distill}}(\mathbf{z}_{i}^{T},\mathbf{z}_{i}^{S})$ (1)
where $y_{i}$ represents the true label for the input $x_{i}$ and $\lambda$ is a hyperparameter that balances the supervised loss and the distillation loss. The supervised loss $\mathcal{L}_{\text{sup}}$ is the task-specific alignment error between the network prediction and annotation. For image classification [53, 63, 59, 15], this is typically cross-entropy loss, while for object detection [46, 9], it includes bounding box regression. The distillation loss $\mathcal{L}_{\text{distill}}$ is the mimic error of the student network towards the teacher network, typically implemented as KL divergence between student and teacher outputs [33].
3.2 Discriminative and Consistent Distillation
We develop an objective function that ensures both discriminative and structurally consistent representations between the teacher’s output $\mathbf{z}_{i}^{T}$ and the student’s output $\mathbf{z}_{i}^{S}$. This objective combines a contrastive loss, which discriminatively aligns representations, with a consistency regularization term that preserves structural relationships in the feature space. The objective function is defined as:
$\mathcal{L}_{\text{kd}}(\mathbf{z}_{i}^{T},\mathbf{z}_{i}^{S})=\mathcal{L}_{\text{contrast}}(\mathbf{z}_{i}^{T},\mathbf{z}_{i}^{S})+\alpha\cdot\mathcal{L}_{\text{consist}}(\mathbf{z}_{i}^{T},\mathbf{z}_{i}^{S})$ (2)
where $\alpha$ is a hyperparameter that balances the contrastive loss $\mathcal{L}_{\text{contrast}}$ for discriminative learning and the consistency regularization term $\mathcal{L}_{\text{consist}}$ for preserving structural relationships.
Discriminative distillation.
In our approach, we employ contrastive learning to align teacher and student representations at the instance level. This process creates similarity between representations of the same input while pushing apart those from different inputs [68]. Through this discriminative mechanism, the student network learns to mirror the teacher’s ability to distinguish between distinct data points.
Instance contrastive learning [70] extends class-wise supervision to its logical extreme by treating each individual instance as its own class. However, this creates a practical challenge: with the number of ”classes” matching the number of training instances, implementing a traditional softmax layer becomes computationally intractable. We resolve this challenge by implementing Noise Contrastive Estimation (NCE) to approximate the softmax, enabling instance-level discrimination without explicit class boundaries:
$\mathcal{L}_{\text{contrast}}(\mathbf{z}_{i}^{T},\mathbf{z}_{i}^{S})=-\log\frac{\exp(\phi(\mathbf{z}_{i}^{S},\mathbf{z}_{i}^{T})/\tau+b)}{\sum_{j=1}^{N}\exp(\phi(\mathbf{z}_{i}^{S},\mathbf{z}_{j}^{T})/\tau+b)}$ (3)
where $\phi(\mathbf{u},\mathbf{v})=\mathbf{u}^{\top}\mathbf{v}/|\mathbf{u}||\mathbf{v}|$ represents the cosine similarity function, with $\tau$ serving as the temperature parameter, $b$ as a bias parameter, and $N$ as the total number of negatives. This formulation effectively transforms into a cross-entropy loss, where each student representation $\mathbf{z}_{i}^{S}$ must identify its corresponding teacher representation $\mathbf{z}_{i}^{T}$ among all other teacher representations in the batch. The objective essentially becomes a classification task: student embeddings must ”classify” their matching teacher embeddings correctly, with the normalized similarities acting as logits and positive pair indices as class labels. The parameters $\tau$ and $b$ provide fine-grained control over this classification process, determining its sharpness and scale respectively.
Consistent distillation.
The consistency loss regularizes the student model to maintain the structural relationships in the teacher model’s representations. Unlike the contrastive term, which operates at the instance level, the consistency loss considers the distributional patterns. The student’s distribution is defined as the similarity between student’s instance $i$ and all other instances $j$ in the batch, processed through a softmax layer:
$p_{i}^{S}(j)=\frac{\exp(\phi(\mathbf{z}_{i}^{S},\mathbf{z}_{j}^{T})/\tau+b)}{\sum_{k=1}^{N}\exp(\phi(\mathbf{z}_{i}^{S},\mathbf{z}_{k}^{T})/\tau+b)}$ (4)
Similarly for the teacher:
$p_{i}^{T}(j)=\frac{\exp(\phi(\mathbf{z}_{i}^{T},\mathbf{z}_{j}^{S})/\tau+b)}{\sum_{k=1}^{N}\exp(\phi(\mathbf{z}_{j}^{T},\mathbf{z}_{k}^{S})/\tau+b)}$ (5)
To achieve this, we adopt a relational consistency approach, which preserves the distributional patterns captured by the teacher model across instances. By aligning the pairwise relationships between instances in the student and teacher embeddings, the model maintains the structural integrity of the teacher’s learned representations. Through KL divergence between the student and teacher similarity distributions, this approach matches not only individual representations but also the spatial configuration of all instances in the embedding space, ensuring robustness and transferability. The consistency regularization term ensures that the student model learns to preserve the structural relationships present in the teacher’s representations by minimizing the KL divergence between these distributions:
$\mathcal{L}_{\text{consist}}(\mathbf{z}_{i}^{T},\mathbf{z}_{i}^{S})=D_{\text{KL}}(\mathbf{p}_{i}^{S}\|\mathbf{p}_{i}^{T})=\sum_{j=1}^{N}p_{i}^{S}(j)\log\frac{p_{i}^{S}(j)}{p_{i}^{T}(j)}$ (6)
where $D_{\text{KL}}$ denotes the KL divergence between the distributions $\mathbf{p}_{i}^{T}$ and $\mathbf{p}_{i}^{S}$, ensuring that the student model maintains similar relational patterns as the teacher model across different inputs.
The combination of contrastive loss and consistency regularization ensures that the learned representations are both discriminative and structurally consistent with the teacher model. This is formalized by the following theorem:
Final objective.
The final objective function, which includes the supervised loss and standard KL divergence, is given by:
$\mathcal{L}=\mathcal{L}_{\text{sup}}(y_{i},\mathbf{z}_{i}^{S})+\lambda\cdot\mathcal{L}_{\text{distill}}(\mathbf{z}_{i}^{T},\mathbf{z}_{i}^{S})+\beta\cdot\mathcal{L}_{\text{kd}}(\mathbf{z}_{i}^{T},\mathbf{z}_{i}^{S})$ (7)
where $\beta$ is a hyperparameter that balances $\mathcal{L}_{\text{kd}}$.
3.3 Implementation Details
We implement the objective using mini-batch stochastic gradient descent. The representations $\mathbf{z}_{i}^{T}$ and $\mathbf{z}_{i}^{S}$ are obtained from the last layer of the teacher and student models, respectively. We further encode $\mathbf{z}_{i}^{T}$ and $\mathbf{z}_{i}^{S}$ using a projection head to match the dimensions. The projection head is trained using stochastic gradient descent as well. This ensures that the representations from both models are compatible for comparison and alignment. Additionally, we normalize the outputs $\mathbf{z}_{i}^{T}$ and $\mathbf{z}_{i}^{S}$ before computing the loss, ensuring that the representations lie on a unit hypersphere. This ensures that the representations from both models are compatible for comparison and alignment.
Memory-efficient sampling.
Instead of using a large memory buffer for contrasting representations as in CRD [66], we use the negative samples that naturally co-exist within the batch. This approach significantly reduces memory requirements while maintaining effective contrastive learning. Our in-batch negative sampling eliminates the need to maintain and update large memory banks that store thousands of feature vectors, which can consume significant GPU memory It also simplifies the implementation by removing the complexity of memory bank management, including challenges related to feature staleness and queue maintenance [30]. Unlike memory banks that may contain stale features from earlier training iterations, our approach always uses the most recent representations within the current batch.
Learnable temperature.
Contrary to contrastive learning objectives that use a constant temperature parameter, we parameterize the temperature using $\exp(\tau)$ where $\tau$ is a learnable parameter, along with a learnable bias $b$. For a batch of normalized embeddings $\mathbf{z}_{i}^{S}$ and $\mathbf{z}_{i}^{T}$, we compute the similarity matrix through:
$\ell_{ij}=\phi(\mathbf{z}_{i}^{S},\mathbf{z}_{j}^{T})\cdot\exp(\tau)+b$ (8)
where $\phi(\mathbf{u},\mathbf{v})=\mathbf{u}^{T}\mathbf{v}/|\mathbf{u}||\mathbf{v}|$ is implemented efficiently as a normalized matrix multiplication. The exponential parameterization ensures the temperature remains positive while allowing unconstrained optimization of $\tau$, which is clamped to $[0,\tau_{\text{max}}]$ for numerical stability. The learnable bias $b$ provides an additive degree of freedom that helps adjust the logit scale. This adaptive approach allows the model to automatically tune the contrast level and logit scaling during training, leading to more robust knowledge transfer compared to fixed hyperparameter approaches.
4 Experiments
We evaluate our DCD framework in the KD task for model compression of a large network to a smaller one, similar to [66]. Beyond classification, we also validate our method’s effectiveness on the more challenging object detection task.
4.1 Experimental Setup
We implement DCD in PyTorch following the implementation of [66].
Datasets.
We conduct experiments on four popular datasets for model compression: (1) CIFAR-100 [40] contains 50,000 training images with 500 images per class and 10,000 test images. (2) ImageNet ILSVRC-2012 [18] includes 1.2 million images from 1,000 classes for training and 50,000 for validation. (3) STL-10 [17] consists of a training set of 5,000 labeled images from 10 classes and 100,000 unlabeled images, and a test set of 8,000 images. (4) Tiny ImageNet [18] comprises 200 classes, each with 500 training images and 50 validation images. (5) MS-COCO [43] contains 118,000 training images and 5,000 validation images with annotations for object detection.
Setup.
We experiment with student-teacher combinations of different capacities, such as ResNet [29] or Wide ResNet (WRN) [75], VGG [64], MobileNet [62], and ShuffleNet [78, 49]. We set $\alpha=0.5$ and $\beta=1$. The hyperparameter $\lambda$ is set to $1.0$ for the KL divergence loss to maintain consistency with [66, 7, 8]. Both the student and teacher outputs are projected to a 128-dimensional space using a projection head consisting of a single linear layer, followed by $\ell_{2}$ normalization. We empirically set $\tau_{\text{max}}=10.0$.
Comparison.
We compare our approach to the following state-of-the-art methods: (1) KD [33]; (2) FitNets [61]; (3) AT [74]; (4) SP [67]; (5) CC [58]; (6) VID [1]; (7) RKD [56]; (8) PKT [57]; (9) AB [32]; (10) FT [37]; (11) FSP [72]; (12) NST [36]; (13) CRD [66]; (14) OFD [31]; (15) WSLD [81]; (16) IPWD [55]; (17) CTKD [42].
Computational cost.
While CRD’s memory bank requires approximately 8MB per class ($16$k features ×$\times 128$ dimensions ×$\times 4$ bytes), DCD’s in-batch sampling needs only 0.13MB total ($256\times 128\times 4$ bytes). This efficiency extends to training time - on a 4-GPU machine, DCD completes ImageNet training in approximately 72 hours compared to CRD’s 88 hours, representing an 18% reduction in training time.
4.2 Main Results
We benchmark our method on image classification and object detection tasks.
Results on CIFAR-100.
Table 1 compares top-1 accuracies of various KD methods on CIFAR-100. Our DCD+KD achieves superior performance, surpassing the teacher network by +0.45% in same-architecture (WRN-40-2 to WRN-16-2) and +0.90% in cross-architecture (WRN-40-2 to ShuffleNet-v1) scenarios. The method shows significant improvements over baseline students: +2.82% for same-architecture and +5.25% for cross-architecture pairs, outperforming memory bank-based CRD. DCD alone performs slightly below CRD, which uses a large $16$k-feature memory bank. However, DCD with KD achieves better results by combining complementary objectives: KD’s soft targets provide direct class-level supervision through logit-space KL divergence, while DCD ensures feature-space consistency.
Table 1: Test top-1 accuracy (%) on CIFAR-100 of student networks trained with various distillation methods across different teacher-student architectures. Architecture abbreviations: W: WideResNet, R: ResNet, MN: MobileNet, SN: ShuffleNet. Results adapted from [66]. Results for our method are averaged over five runs.
Same architecture Different architecture
Teacher W-40-2 W-40-2 R-56 R-110 R-110 R-32x4 VGG-13 VGG-13 R-50 R-50 R-32x4 R-32x4 W-40-2
Student W-16-2 W-40-1 R-20 R-20 R-32 R-8x4 VGG-8 MN-v2 MN-v2 VGG-8 SN-v1 SN-v2 SN-v1
Teacher 75.61 75.61 72.34 74.31 74.31 79.42 74.64 74.64 79.34 79.34 79.42 79.42 75.61
Student 73.26 71.98 69.06 69.06 71.14 72.50 70.36 64.60 64.60 70.36 70.50 71.82 70.50
KD [33] 74.92 73.54 70.66 70.67 73.08 73.33 72.98 67.37 67.35 73.81 74.07 74.45 74.83
FitNet [61] 73.58 72.24 69.21 68.99 71.06 73.50 71.02 64.14 63.16 70.69 73.59 73.54 73.73
AT [74] 74.08 72.77 70.55 70.22 72.31 73.44 71.43 59.40 58.58 71.84 71.73 72.73 73.32
SP [67] 73.83 72.43 69.67 70.04 72.69 72.94 72.68 66.30 68.08 73.34 73.48 74.56 74.52
CC [58] 73.56 72.21 69.63 69.48 71.48 72.97 70.81 64.86 65.43 70.25 71.14 71.29 71.38
VID [1] 74.11 73.30 70.38 70.16 72.61 73.09 71.23 65.56 67.57 70.30 73.38 73.40 73.61
RKD [56] 73.35 72.22 69.61 69.25 71.82 71.90 71.48 64.52 64.43 71.50 72.28 73.21 72.21
PKT [57] 74.54 73.45 70.34 70.25 72.61 73.64 72.88 67.13 66.52 73.01 74.10 74.69 73.89
AB [32] 72.50 72.38 69.47 69.53 70.98 73.17 70.94 66.06 67.20 70.65 73.55 74.31 73.34
FT [37] 73.25 71.59 69.84 70.22 72.37 72.86 70.58 61.78 60.99 70.29 71.75 72.50 72.03
FSP [72] 72.91 n/a 69.95 70.11 71.89 72.62 70.33 58.16 64.96 71.28 74.12 74.68 76.09
CRD [66] 75.48 74.14 71.16 71.46 73.48 75.51 73.94 69.73 69.11 74.30 75.11 75.65 76.05
CRD+KD [66] 75.64 74.38 71.63 71.56 73.75 75.46 74.29 69.94 69.54 74.58 75.12 76.05 76.27
OFD [31] 75.24 74.33 70.38 n/a 73.23 74.95 73.95 69.48 69.04 n/a 75.98 76.82 75.8
WSLD [81] n/a 73.74 71.53 n/a 73.36 74.79 n/a n/a 68.79 73.80 75.09 n/a 75.23
IPWD [55] n/a 74.64 71.32 n/a 73.91 76.03 n/a n/a 70.25 74.97 76.03 n/a 76.44
CTKD [42] 75.45 73.93 71.19 70.99 73.52 n/a 73.52 68.46 68.47 n/a 74.78 75.31 75.78
DCD (ours) 74.99 73.69 71.18 71.00 73.12 74.23 73.22 68.35 67.39 73.85 74.26 75.26 74.98
DCD+KD (ours) 76.06 74.76 71.81 72.03 73.62 75.09 73.95 69.77 70.03 74.08 76.01 76.95 76.51
Transferability of representations.
Table 2 compares the top-1 test accuracy of WRN-16-2 (student) distilled from WRN-40-2 (teacher) and evaluated on STL-10 and Tiny ImageNet. The student is trained on CIFAR-100, either directly or via distillation, and serves as a frozen feature extractor with a linear classifier. We assess how well different distillation methods enhance feature transferability. Our results show that DCD, both standalone and combined with KD, significantly improves transferability.
Table 2: Test top-1 accuracy (%) of WRN-16-2 (student) distilled from WRN-40-2 (teacher). In this setup, the representations learned from the CIFAR-100 dataset are transferred to the STL-10 and Tiny ImageNet datasets.
Teacher Student KD [33] AT [74] FitNet [61] CRD [66] CRD+KD [66] DCD DCD+KD
[…]
Results on ImageNet.
Table 3 presents the top-1 accuracy of student networks trained with various distillation methods on ImageNet. The results highlight the effectiveness of our approach in large-scale settings, demonstrating its ability to distill knowledge from complex models and enhance student performance. Our method consistently surpasses KD and achieves competitive results across different architectures.
Results on COCO.
Table 4 shows our performance on the MS-COCO object detection task. Following [79], we adopt Faster R-CNN [60] with Feature Pyramid Network (FPN) [44] as our detection framework. We evaluate two teacher-student scenarios: ResNet-101 to ResNet-50 and ResNet-50 to MobileNet-V2. For ResNet-101 to ResNet-50 distillation, DCD+KD achieves 39.01 AP and 60.07 AP50, surpassing the baseline by +1.08 AP. When distilling from ResNet-50 to MobileNet-v2, we obtain substantial improvements of +3.81 AP over the baseline, demonstrating effective knowledge transfer even across different architectures.
4.3 Visualization
Table 3: Test top-1 accuracy (%) on ImageNet validation set for student networks trained with various distillation methods across different teacher-student architectures. Results for our method are based on a single run.
Teacher Student KD [33] AT [74] SP [67] CC [58] RKD [56] CRD [66] DCD DCD+KD
[…]
Table 4: Object detection performance on MS-COCO val2017 using Faster R-CNN with FPN backbone. We use mean Average Precision (AP) and AP at IoU thresholds of 0.5 and 0.75 (AP50, AP75). Results for our method are based on a single run.
Method ResNet-101 $\rightarrow$ ResNet-18 ResNet-101 $\rightarrow$ ResNet-50 ResNet-50 $\rightarrow$ MobileNet-v2
AP AP50 AP75 AP AP50 AP75 AP AP50 AP75
Teacher 42.04 62.48 45.88 42.04 62.48 45.88 40.22 61.02 43.81
Student 33.26 53.61 35.26 37.93 58.84 41.05 29.47 48.87 30.90
KD [33] 33.97 54.66 36.62 38.35 59.41 41.71 30.13 50.28 31.35
FitNet [74] 34.13 54.16 36.71 38.76 59.62 41.80 30.20 49.80 31.69
DCD (ours) 36.98 57.44 39.79 39.01 60.07 42.88 33.28 52.97 35.15
(a) Vanilla
Mean: 0.24, Max: 1.66
(b) KD [33]
Mean: 0.09, Max: 0.49
(c) CRD [66]
Mean: 0.23, Max: 1.56
(d) CRD+KD [66]
Mean: 0.10, Max: 0.57
(e) DCD (ours)
Mean: 0.22, Max: 1.59
(f) DCD+KD (ours)
Mean: 0.09, Max: 0.51
Figure 2: Correlation matrix of the average logit difference between teacher and student logits on CIFAR-100. We use WRN-40-2 as the teacher and WRN-40-1 as the student. Methods have been re-implemented according to [66].
(a) Teacher
(b) Vanilla
(c) KD [33]
(d) CRD [66]
(e) CRD+KD [66]
(f) DCD (ours)
(g) DCD+KD (ours)
Figure 3: t-SNE visualizations of embeddings from teacher and student networks on CIFAR-100 (first 20 classes). We use WRN-40-2 as the teacher and WRN-40-1 as the student. Methods have been re-implemented according to [66].
We present comprehensive visualizations that analyze the learned representations and knowledge transfer patterns across different distillation approaches, providing qualitative insights.
Inter-class correlations.
Figure 2 evaluates the effectiveness of distillation methods on the CIFAR-100 KD task using WRN-40-2 (teacher) and WRN-40-1 (student). We compare models trained without distillation, with KL divergence [33], with CRD [66], and our proposed DCD method. DCD achieves strong correlation alignment between teacher and student logits, reducing discrepancies in their correlation matrices.
t-SNE visualization.
Figure 3 presents t-SNE [69] visualizations of embeddings from WRN-40-2 (teacher) and WRN-40-1 (student) on CIFAR-100. Compared to standard training and [74], DCD better aligns student embeddings with the teacher, preserving structural relationships in the feature space. The preservation of semantic relationships between class clusters demonstrates our method’s effectiveness in transferring the teacher’s knowledge organization.
4.4 Ablation Study
We analyze our method through ablation experiments presented in Figures 4 and 5. We first study the effects of consistency regularization and temperature scaling, then investigate the sensitivity to hyperparameters $\alpha$, $\beta$, and $\lambda$.
(a) DCD variants without KD
(b) DCD variants with KD
Figure 4: Ablation study results on CIFAR-100. We show results for discriminative training ($\alpha=0$), discriminative and consistent training ($\alpha=0.5$, $\tau=0.07$, $b=0$), and our proposed DCD approach ($\alpha=0.5$, trainable $\tau$ and $b$). The colors correspond to each respective variant. (a) compares DCD variants without knowledge distillation, while (b) shows improvements when combined with KD. Results are based on a single run.
(a) Effect of varying $\alpha$
(b) Effect of varying $\beta$
(c) Effect of varying $\lambda$
Figure 5: Ablation study results on CIFAR-100 using WRN-40-2 as the teacher and WRN-16-2 as the student. (a) Effect of the internal DCD coefficient $\alpha$ in Equation 2. (b) Effect of DCD loss coefficient $\beta$ in Equation 7. (c) Effect of loss coefficient $\lambda$ in Equation 7. Results are averaged over five runs.
Ablation on DCD.
To examine the effectiveness of consistency regularization and temperature scaling, we conduct comprehensive ablation studies on CIFAR-100, as shown in Figure 4. Starting with a pure discriminative variant ($\alpha=0$), we observe that adding consistency regularization with fixed temperature ($\alpha=0.5$, $\tau=0.07$, $b=0$) improves performance across all architectures. Our proposed method with trainable temperature parameters further enhances the results, achieving improvements of up to +1.69% without KD and +2.21% with KD over the baseline. These results demonstrate that both consistency regularization and adaptive temperature scaling contribute significantly to the method’s performance.
Ablation on $\bm{\alpha}$.
We then tested different values for the loss coefficient $\alpha$ of Equation 2: $\alpha=\{0.01,0.1,0.3,0.5,0.7,1,2,5\}$ while setting $\beta=1$ and $\lambda=0$. For all the following ablations we use WRN-40-2 as teacher and WRN-16-2 as student on CIFAR-100. As shown in Figure 5(a), our method remains robust across changes in $\alpha$, with no significant difference in performance at low or high values. This robustness can be attributed to the adaptive temperature scaling mechanism, which enables automatic tuning of contrast level and logit scaling during training, providing more stable knowledge transfer.
Ablation on $\bm{\beta}$.
We varied $\beta$ of Equation 7 from $\beta=0.1$ to $\beta=100$, considering values of $\beta=\{0.1,0.5,1,2,5,10,50,100\}$ while fixing $\alpha=0.5$ and $\lambda=0$. As illustrated in Figure 5(b), extremely high $\beta$ values cause significant degradation in performance due to the overwhelming contribution of the DCD loss relative to other loss terms. Very low values of $\beta$ also lead to a slight decrease in performance. The optimal range for $\beta$ is between $\beta=0.5$ and $\beta=10$, suggesting that the DCD loss should be weighted similarly to other loss terms for the best results.
Ablation on $\bm{\lambda}$.
While $\lambda$ of Equation 7 is typically set to $\lambda=1.0$ [66, 7, 8], we tested values from $\lambda=0.1$ to $\lambda=10$, considering values of $\lambda=\{0.1,0.5,1,2,5,10\}$. For these experiments we fixed $\alpha=0.5$ and $\beta=1$. As shown in Figure 5(c), performance remains stable across all tested values, with the best results achieved at $\lambda=1.0$ (we also found that higher values of $\lambda=50$ and $\lambda=100$ lead to training collapse, not shown in figure). This confirms the common practice of setting $\lambda=1.0$ in prior work is optimal.
5 Conclusions
We have presented DCD, a knowledge distillation method that combines contrastive learning with consistency regularization to improve the traditional KD process. Our method achieves state-of-the-art performance through memory-efficient in-batch negative sampling and adaptive temperature scaling, eliminating the need for large memory banks while automatically tuning contrast levels during training. Through extensive experimentation across CIFAR-100, ImageNet, COCO, STL-10, and Tiny ImageNet datasets, we have demonstrated significant improvements over existing methods. Unlike previous methods such as CRD [66] that require large memory banks or WSLD [81] and IPWD [55] that focus solely on instance discrimination, our approach achieves superior performance while being more memory-efficient and capturing both local and global structural information. The effectiveness of DCD has been validated in both same-architecture and cross-architecture scenarios, with student models in several cases exceeding their teachers’ performance. We hope this work will inspire future research in knowledge distillation.
Acknowledgments
We would like to express our gratitude to Andreas Floros for his valuable feedback, particularly his assistance with equations, notations, and insightful discussions that greatly contributed to this work. We also acknowledge the computational resources and support provided by the Imperial College Research Computing Service (http://doi.org/10.14469/hpc/2232), which enabled our experiments.